OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 5x - y = 27 5x + 4y = 42

OpenStudy (chaise):

5x - y = 27 5x + 4y = 42 Take them away from eachother to eliminate the x -5y = -15 y = 3 5x - y = 27 5x -3 = 27 5x = 30 x =6 (x,y) (6,3) I've done two for you - I'm not doing your entire homework